# Sandwiched singularities and nearly Lefschetz fibrations

### v2

@@ 4.2. Braided wiring diagrams:
By definition, the monodromy over $\alpha|_{[0,t_{1}-\delta]}$ is $b_{0}$ and the monodromy over $\alpha|_{[t_{i}+\delta,t_{i+1}-\delta]}$ is $b_{i}$ for $i=1,\dots,N-1$. Through the local models for the intersection of a pencil of transverse complex lines and the tangency, we see that the monodromy over $h_{i}$ is $\phi_{i}$ as defined in the statement of the lemma, and the monodromy over $g_{i}$ is $\Delta_{j,j+k}^{2}$ at an intersection singularity or free marked point and $\Delta_{j,j+1}$ at a tangency. Let $\tau_{C}$ denote a positive Dehn twist about a curve $C$ and let $t_{\gamma}$ denote a boundary interchange about an arc $\gamma$. Then the monodromy $m_{l_{i}}$ about $l_{i}$ is given by
@@
Therefore $m_{l_{i}}=\tau_{V_{i}}$ or $t_{V_{i}}$ where $V_{i}$ is the curve or arc given in the statement of the lemma.
@@
Let $P$ be the fiber of the planar spinal open book, identified with a disk with $r$ holes of equal radius centered on the real axis and labeled $1,\dots,r$ in order of increasing real part. Note that there is a surjective map from the group of diffeomorphisms of $P$ to the braid group on $r$ strands $\iota:\operatorname{Diff}(F)\to B_{r}$ induced from the inclusion of $P$ into the punctured plane whose kernel is generated by Dehn twists about curves parallel to boundary components of $P$.
If $V_{1}$ is a vanishing cycle (closed curve), choose a diffeomorphism $\psi_{0}:P\to P$ such that $\psi_{0}(V_{1})$ is convex. Let $b_{0}=\iota(\psi_{0})$. Let $A_{1}=\psi_{0}(V_{1})$, $S_{1}$ the intersection singularity of the strands corresponding to holes enclosed by $A_{1}$ which we will mark with a point $p_{1}$, and $\phi_{1}$ the half-twist on the strands corresponding to holes enclosed by $A_{1}$. Note that if $A_{1}$ only encloses a single hole, $S_{1}$ is not a singularity, but we will still mark it with a point $p_{1}$. If $V_{1}$ is a vanishing arc, choose a diffeomorphism $\psi_{0}:P\to P$ such that $\psi_{0}(V_{1})$ lies on the real axis of the disk and connects two adjacent boundary components, holes $j$ and $j+1$. Then let $b_{0}=\iota(\psi_{0})$, $A_{1}=\psi_{0}(V_{1})$, $S_{1}$ the tangency between strands $j$ and $j+1$, and $\phi_{1}=id$.
Inductively, we will define $b_{k-1}$, $A_{k}$ and $\phi_{k}$ from previously defined values of $b_{i-1}$, $A_{i}$, $\phi_{i}$ for $i<k$. Choose a diffeomorphism $\psi_{k-1}:P\to P$ such that $A_{k}:=\phi_{k-1}\circ b_{k-2}\circ\cdots\circ b_{1}\circ\phi_{1}\circ b_{0}(V_{k})$ is a convex curve or straight line arc on the real axis. Let $b_{k-1}=\iota(\psi_{k-1})$. If $A_{k}$ is a closed curve, let $S_{k}$ be the intersection singularity of the strands enclosed by $A_{k}$, marked with a point $p_{k}$, and $\phi_{k}$ the half twist on the strands enclosed by $A_{k}$. Again, if $A_{k}$ happens to only enclose a single hole, $S_{k}$ will not be a singularity, but we will still mark the point $p_{k}$. If $A_{k}$ is an arc, let $S_{k}$ be the tangency between the two adjacent strands connected by $A_{k}$, and let $\phi_{k}=id$. We let $b_{N}=id$.
@@ 5.1. Detecting the incidence matrix
In [PS23], we described the boundary data in terms of a fixed open book rather than a marking. Because of Lemmas 3.11 and 3.10, fixing the marking is equivalent to fixing the isotopy class of the spinal open book, if one also knows the multiplicities of the curvettas $C_{i}$ at 0. Indeed, by Lemma 3.10, the marking gives binding components $b_{i}$ for $i=1,\dots,m$ that correspond to the curvettas $C_{i}$. The remaining outer binding component $b_{0}$ is then uniquely determined as well: the auxiliary (1)(-1) vertices adjacent to $E(C_{i})$ determine the blow downs of the augmented resolution graph, and therefore the vertex $E_{0}$.
@@ 5.2. Unexpected fillings
Figure 13. Above: The sandwiched dual graph $G$ from Figure 1. Below: A sandwiched dual graph $K$ containing $G$. Unlabeled small vertices have self-intersection 2-2. The link of the sandwiched singularity with normal crossing resolution $K$ admits at least $5$ pairwise non-homeomorphic unexpected Stein fillings. As in the proof of Theorem 1.2 with $N=4$, $K$ is obtained from $G$ by adding a vertex $v^{*}$ to the left-most 3-3 vertex associated with $E_{0}$ such that the self-intersection on $v^{*}$ is 15-15, with $13$ additional legs attached to $v^{*}$ of length $14$ with all 2-2-vertices, and 22 additional legs, each of length $26$ and with all 2-2-vertices, attached to the 3-3 vertex associated with $E(C_{1})=E(C_{2})$ on the far right.
@@
A similar statement is true for the general case, but one can only turn the open book inside out if there is another boundary compoment of multiplicity 1. We leave the details to the reader.
We can use the previous remark to generate a new kind of a non-realizable arrangement. All of our unexpected fillings above are using the same type of examples as a key ingredient: we need a collection of lines with distinct slopes to be a subgerm of the given decorated germ. By contrast, Proposition 5.9 gives examples where all the curvettas have a high-order tangency with one another. Although the construction is based on a trick from Remark 5.8 and doesn’t produce any new unexpected Stein fillings, perhaps these examples shed some more light on arrangements combinatorially non-realizable by picture deformations. It would be interesting to have further analytic criteria to detect such arrangements, beyond [PS23, Lemma 7.5].
@@
It is easy to see how $\Gamma^{\prime}$ is related to $\Gamma\cup l$ and $\Gamma$. In the new open book factorization, the Dehn twists corresponding to intersection points between components of $\Gamma_{i}$ remain unchanged and are in the obvious 1-1 correspondence with the Dehn twists in the old factorization. What changes is the Dehn twists involving the extra hole: a Dehn twist enclosing $h_{l}$ and $h_{i}$ now becomes a Dehn twist in the new open book that encloses $h_{o}$ and all of the $h_{j}$’s except $h_{i}$. By our assumptions, these Dehn twists are precomposed with the rest of the monodromy, and the arrangement $\Gamma^{\prime}$ is build by adding a disk $\Gamma_{0}$ to $\Gamma$ and precomposing with extra wiring with transverse intersection points of $m$-tuples $(\Gamma_{0},\Gamma_{1},\dots,\hat{\Gamma}_{i},\dots,\Gamma_{m})$, in order. (Here, the hat indicates that the corresponding disk is not included.) It is easiest to visualize this wiring diagram by placing the intersection points of $\Gamma_{i}$ with the $\mathbb{C}$-fiber of the diagram at roots of unity for $i=1,\dots,m$, with $\Gamma_{0}$ at the origin, and making all strands except one meet at the origin, in order.

### v3

@@ 4.2. Braided wiring diagrams:
We will compute the monodromy over the loops $l_{i}$ as the composition of the monodromy over the elementary open paths that we decomposed $l_{i}$ into. The monodromy over each of these open paths is well-defined because we can identify different fibers using the projection from $\mathbb{C}^{2}$ to $\mathbb{C}$ onto the second coordinate, and we will assume (by an isotopy if needed) that the surface intersects the fibers at the start and end of these paths in the same set of points under this identification. The corresponding elementary monodromy is determined via local models in each case. By definition, the monodromy over $\alpha|_{[0,t_{1}-\delta]}$ is $b_{0}$ and the monodromy over $\alpha|_{[t_{i}+\delta,t_{i+1}-\delta]}$ is $b_{i}$ for $i=1,\dots,N-1$. As in the statement of the lemma, we will fix arbitrary lifts of these braids to the mapping class group of the fiber $P$, keeping the same notation. From a local model for the intersection of a pencil of transverse complex lines, we see that the braid monodromy over $g_{i}$ is the full twist $\Delta_{j,j+k}^{2}$ on the corresponding strands. For the Lefschetz fibration on the complement of the tubular neighborhoods of the strict transforms of the curves in the blowup, the local model as in [PS23, Lemma 3.2] gives the the monodromy over $g_{i}$ as the positive Dehn twist $\tau_{A_{i}}$ along the convex simple closed curve $A_{i}$ enclosing the corresponding holes. For the free marked point, we get a positive Dehn twist around a single hole. In the case of a tangency, the local model gives the braid monodromy over $g_{i}$ as the half twist $\Delta_{j,j+1}$. The monodromy of the nearly Lefschetz fibration is the positive boundary interchange $t_{A_{j}}$ on the arc $A_{j}$, by the local model as in Figure 4. As explained before the statement of the lemma, the local models for the fibrations determine the boundary parallel Dehn twists in the monodromy, in addition to the information read off from the braid monodromy of the arrangement.
Similarly, the local model for the intersection of complex lines tells us that the braid monodromy over the path $h_{i}$ is the negative half twist $\phi_{j}=\Delta_{i,i+k}^{-1}$ on the corresponding strands, as defined in the statement of the lemma. (The negative sign is due to the clockwise orientation of $h_{i}$.) In the case of tangency between strands $i$ and $i+1$, the braid monodromy over the path $h_{i}$ is the negative half twist $\phi_{j}=\sigma_{i}^{-1}$: we use the local model and pay attention to the path orientation and the fact that it goes “in the back” of the diagram. For the nearly Lefschetz fibration on the complement of the tubular neighborhoods of the strict transforms of the curves in the blowup, we can determine the boundary Dehn twists from the local models. However, we will see that the maps $\phi_{j}$ only enter into the total monodromy expression as conjugations. The boundary Dehn twists will not affect the generalized Hurwitz equivalence class of the fibration, therefore, we can choose to ignore them at this step.
As above, let $\tau_{C}$ denote a positive Dehn twist about a curve $C$ and let $t_{\gamma}$ denote a boundary interchange about an arc $\gamma$. Then the monodromy $m_{l_{i}}$ of the nearly Lefschetz fibration $W_{(\Gamma,p)}$ about $l_{i}$ is given by
@@
Therefore $m_{l_{i}}=\tau_{V_{i}}$ or $t_{V_{i}}$ where $V_{i}$ is the curve or arc given in the statement of the lemma. Our constructions had an arbitrary choice of lifts of braid elements $b_{i}$ and $\phi_{i}$ to the mapping class group of the disk with holes, with an undeterminacy coming from boundary parallel Dehn twists. In the formulas we obtained, the boundary parallel Dehn twists commute with the Dehn twists along the vanishing cycles, so the vanishing cycles are not affected by this undeterminacy. Vanishing arcs are changed by the framing conjugation move: a boundary interchange $t_{\gamma}$ is replaced by $\tau^{-1}_{\delta}t_{\gamma}\tau_{\delta}$, conjugating by the Dehn twist around a boundary parallel curve. This gives the same generalized Hurwitz equivalence class, [BH16a], [MRW24, Definition 2.14]. ∎
@@
Let $P$ be the fiber of the planar spinal open book, identified with a disk with $r$ holes of equal radius centered on the real axis and labeled $1,\dots,r$ in order of increasing real part. As above, there is a surjective map $\iota:Map(P)\to B_{r}$ from the mapping class group of $P$ to the braid group, with kernel generated by the boundary parallel Dehn twists.
If $V_{1}$ is a vanishing cycle (closed curve), choose a diffeomorphism $\psi_{0}:P\to P$ such that $\psi_{0}(V_{1})$ is convex and encloses holes $i,\dots,i+k$. Let $b_{0}:=\iota(\psi_{0})$. Let $A_{1}:=\psi_{0}(V_{1})$, $S_{1}:=I^{i}_{i+k}$ marked with a point $p_{1}$, and $\phi_{1}:=\Delta_{i,i+k}^{-1}$. Note that if $A_{1}$ only encloses a single hole, $S_{1}$ is not a singularity, but we will still mark it with a point $p_{1}$. If $V_{1}$ is a vanishing arc, choose a diffeomorphism $\psi_{0}:P\to P$ such that $\psi_{0}(V_{1})$ lies on the real axis of the disk and connects two adjacent boundary components, holes $i$ and $i+1$. Then let $b_{0}:=\iota(\psi_{0})$, $A_{1}:=\psi_{0}(V_{1})$, $S_{1}:=T_{i}$, and $\phi_{1}:=\sigma_{i}^{-1}$.
Inductively, we will define $b_{l-1}$, $A_{l}$ and $\phi_{l}$ from previously defined values of $b_{j-1}$, $A_{j}$, $\phi_{j}$ for $j<l$. Choose a diffeomorphism $\psi_{l-1}:P\to P$ such that $A_{l}:=\phi_{l-1}\circ b_{l-2}\circ\cdots\circ b_{1}\circ\phi_{1}\circ b_{0}(V_{l})$ is a convex curve or straight line arc on the real axis. Let $b_{l-1}=\iota(\psi_{l-1})$. If $A_{l}$ is a closed curve containing holes $i,\dots,i+k$, let $S_{l}:=I^{i}_{i+k}$, marked with a point $p_{l}$, and $\phi_{l}:=\Delta_{i,i+k}^{-1}$. Again, if $A_{l}$ happens to only enclose a single hole, $S_{l}$ will not be a singularity, but we will still mark the point $p_{l}$. If $A_{l}$ is an arc connecting the $i$ and $i+1$ strands, let $S_{l}:=T_{i}$, and let $\phi_{k}:=\sigma_{i}^{-1}$. We let $b_{N}=id$.
@@ 5.1. Detecting the incidence matrix
In [PS23], we described the boundary data in terms of a fixed open book rather than a marking. Because of Lemmas 3.11 and 3.10, fixing the marking is equivalent to fixing the isotopy class of the spinal open book, if one also knows the multiplicities of the curvettas $C_{i}$ at 0. Indeed, by Lemma 3.10, the marking gives binding components $\beta_{i}$ for $i=1,\dots,m$ that correspond to the curvettas $C_{i}$. The remaining outer binding component $\beta_{0}$ is then uniquely determined as well: the auxiliary (1)(-1) vertices adjacent to $E(C_{i})$ determine the blow downs of the augmented resolution graph, and therefore the vertex $E_{0}$.
@@ 5.2. Unexpected fillings
Figure 13. Above: The sandwiched dual graph $G$ from Figure 1. Below: A sandwiched dual graph $K$ containing $G$. Unlabeled small vertices have self-intersection 2-2. The link of the sandwiched singularity with normal crossing resolution $K$ admits at least $5$ pairwise non-homeomorphic unexpected Stein fillings. As in the proof of Theorem 1.2 with $N=4$, $K$ is obtained from $G$ by adding a vertex $v^{*}$ to the left-most 3-3 vertex associated with $E_{0}$ such that the self-intersection on $v^{*}$ is 15-15, with $13$ additional legs attached to $v^{*}$ of length $14$ with all 2-2-vertices, and $2$ additional legs, each of length $26$ and with all 2-2-vertices, attached to the 3-3 vertex associated with $E(C_{1})=E(C_{2})$ on the far right.
@@
A similar statement is true for the general spinal case, but one can only turn the open book inside out if there is another boundary compoment of multiplicity 1. We leave the details to the reader.
We can use the previous remark to generate a new kind of a non-realizable arrangement. All of our unexpected fillings above are using the same type of examples as a key ingredient: we need a collection of lines with distinct slopes to be a subgerm of the given decorated germ. By contrast, Proposition 5.9 gives examples where all the curvettas have a high-order tangency with one another. The construction is based on a trick from Remark 5.8 and doesn’t produce any new unexpected Stein fillings, these examples shed some more light on arrangements combinatorially non-realizable by picture deformations. We give both a topological proof of Proposition 5.9 and an analytic explanation of same fact in Proposition 5.10.
@@
It is easy to see how $\Gamma^{\prime}$ is related to $\Gamma\cup l$ and $\Gamma$. In the new open book factorization, the Dehn twists corresponding to intersection points between components of $\Gamma_{i}$ remain unchanged and are in the obvious 1-1 correspondence with the Dehn twists in the old factorization. What changes is the Dehn twists involving the extra hole: a Dehn twist enclosing $h_{l}$ and $h_{i}$ now becomes a Dehn twist in the new open book that encloses $h_{o}$ and all of the $h_{j}$’s except $h_{i}$. By our assumptions, these Dehn twists are precomposed with the rest of the monodromy, and the arrangement $\Gamma^{\prime}$ is build by adding a disk $\Gamma_{0}$ to $\Gamma$ and precomposing with extra wiring with transverse intersection points of $m$-tuples $(\Gamma_{0},\Gamma_{1},\dots,\hat{\Gamma}_{i},\dots,\Gamma_{m})$, in order. (Here, the hat indicates that the corresponding disk is not included.) It is easiest to visualize this wiring diagram by placing the intersection points of $\Gamma_{i}$ with the $\mathbb{C}$-fiber of the diagram at roots of unity for $i=1,\dots,m$, with $\Gamma_{0}$ at the origin, and making all strands except one meet at the origin, in order. See Figure 14.
@@
The inside-out topological argument above shows that the arrangement $\Gamma^{\prime}$ is not realizable by a small analytic deformation if the arrangement $\Gamma$ is not. We are able to give a direct analytic proof of this fact. We will assume that the components of $\Gamma$ are smoothly embedded graphical disks (equivalently, the compatible germ $C$ has smooth components). Then a small deformation of this germ also has smooth graphical components. The weights and marked points do not play a role here, and we omit them from the statement below. (The incidence matrix records intersections between the components of $\Gamma$.) Note that the proposition below is stronger than what the topological argument implies: we only need a combinatorial assumption on the incidences of $\Gamma^{\prime}$, without any specific requirements on the wiring diagram.
Proposition 5.10.
Suppose that $\Gamma$ a smooth disk arrangement compatible with a germ $(C,w)$ with smooth components, such that the incidence matrix $\mathcal{I}(\Gamma)$ is not realizable by a small analytic deformation. Consider an arrangement $\Gamma^{\prime}$ related to $\Gamma$ as in Figure 14: $\Gamma^{\prime}$ has an extra disk $\Gamma_{0}$ in addition to the components of $\Gamma$, and additional transverse intersection points of $m$-tuples $(\Gamma_{0},\Gamma_{1},\dots,\hat{\Gamma}_{i},\dots,\Gamma_{m})$. Then the incidence matrix of the arrangement $\Gamma^{\prime}$ is not realizable by a small analytic deformation of a compatible germ.
Proof.
Assume the contrary, namely, there are equations of the form $y=f_{j}(x,s)$ with analytic functions $f_{j}(x,s)$ that encode a disk arrangement with the incidences of $\Gamma^{\prime}$ for $s>0$, specializing to a compatible curve germ for $s=0$. Substracting the equation $y=f_{0}(x,s)$ of the disk $\Gamma_{0}$ from each of the equations, we can assume that $\Gamma_{0}$ is given by $y=0$. Then for each $s>0$, the $m$-tuple intersections $(\Gamma_{0},\Gamma_{1},\dots,\hat{\Gamma}_{i},\dots,\Gamma_{m})$ are points $(x=r_{i}(s),y=0)$ for some analytic functions $r_{i}(s)$, and the components $\Gamma^{\prime}_{1},\dots,\Gamma_{m}^{\prime}$ are given by equations of the form $y=q_{i}(x,s)$, where
$q_{i}(x,s)=(x-r_{1}(s))\dots(x-r_{i-1}(s))(x-r_{i+1}(s))\dots(x-r_{m}(s))g_{i}(x,s).$
Here, $g_{i}(x,s)$ are analytic functions that are non-vanishing for all $(x,s)$: by assumption, the only places where $y=q_{i}(x,s)$ intersects $y=0$ is at $r_{1}(s),\cdots,r_{m}(s)$, and these zeros have multiplicity one because the intersections with $y=0$ are transverse.
Let
$p(x,s)=(x-r_{1}(s))(x-r_{2}(s))\dots(x-r_{m}(s)),$
and consider, for $s>0$, an arrangement given by equations
$y=\frac{p(x,s)}{q_{i}(x,s)}=\frac{(x-r_{i}(s))}{g_{i}(x,s)}.$
For $s=0$, these equations specialize to equations of a reducible curve germ at 0. For $s>0$, the incidences of this new arrangement are the same as the incidences of $\Gamma$: indeed, the above formulas produce the incidences of the family $y=q_{i}(x,s)$ with all the points $(r_{i}(s),0)$ removed. In both cases, incidences between the $i^{th}$ and $j^{th}$ components correspond exactly to $(x,s)$ solutions to $(x-r_{i}(s))g_{j}(x,s)=(x-r_{j}(s))g_{i}(x,s)$. It follows that if the combinatorial arrangement $\Gamma^{\prime}$ admitted a realization by a small analytic deformation, the same would be true for $\Gamma$, a contradiction. ∎
Since we work directly with equations of the curves, the above proposition and its proof can be restated for deformation realizations of arrangements up to isotopy (rather than in the combinatorial sense).
Remark 5.11.
Although it doesn’t seem to lead to interesting topological consequences, there is a more obvious way to create deformation nonrealizable arrangements compatible with germs with smooth components tangent to a high order. We can simply use a known unrealizable arrangement and add a number of points where all the components intersect. Indeed, let $\Gamma$ be an arrangement with $m$ smooth components as in Proposition 5.9. Consider an arrangement $\Gamma^{\prime}$, also with $m$ smooth components and singularities given by transverse multipoints only, such that the incidence relations of $\Gamma^{\prime}$ are given by those of $\Gamma$, plus $k$ additional $m$-points. Suppose that $\Gamma^{\prime}$ can be obtained by a small analytic deformation. As before, we can assume that one of the components of $\Gamma^{\prime}$ is given by the line $y=0$ at all times. If $(x=r_{i}(s),y=0)$ are the $m$-points for $i=1,\dots,k$, then the deformation for the curves producing $\Gamma^{\prime}$ is given by equations of the form
$y=(x-r_{1}(s))\dots(x-r_{k}(s))g_{i}(x,s).$
Here, $g_{i}(x,s)$ are analytic functions (the one corresponding to the curve $y=0$ is identically 0). It is easy to see that functions $g_{i}(x,s)$ give a germ deformation with incidences of $\Gamma$ for $s>0$, so if the incidence matrix of $\Gamma$ is not realizable by a small analytic deformation, neither is the incidence matrix of $\Gamma^{\prime}$.
Using the last two constructions together with [PS23, Lemma 7.5] producing unexpected pseudoline configurations, we can create various non-realizable arrangements with smooth components. However, more general criteria and similar direct analytic proofs of deformation nonrealizability remain elusive. We have no results for arrangements corresponding to cuspidal germs, in particular, no direct explanation for nonrealizability of arrangements of nodal cubics in [BPS25]. It would be very interesting to develop further tools to detect interesting arrangements.